### v1

# Entropy Heat-Mapping: Localizing GPT-Based OCR Errors with Sliding-Window Shannon Analysis

Abstract
Vision–language models such as OpenAI GPT-4o can transcribe mathematical documents directly from images, yet their token-level confidence signals are seldom used to pinpoint local recognition mistakes. We present an entropy-heat-mapping proof-of-concept that turns per-token Shannon entropy into a visual “uncertainty landscape.” By scanning the entropy sequence with a fixed-length sliding window, we obtain hotspots that are likely to contain OCR errors such as missing symbols, mismatched braces, or garbled prose. Using a small, curated set of scanned research pages rendered at several resolutions, we compare the highlighted hotspots with the actual transcription errors produced by GPT-4o. Our analysis shows that the vast majority of true errors are indeed concentrated inside the high-entropy regions. This study demonstrates—in a minimally engineered setting—that sliding-window entropy can serve as a practical, lightweight aid for post-editing GPT-based OCR. All code, sample data, and annotation guidelines are released to encourage replication and further research.
1 Background and Related Work
This section provides context on (i) the information–theoretic lens for understanding OCR, (ii) the extension of GPT from purely text-based to vision–language modeling, (iii) prior efforts leveraging per-token probability distributions for confidence checks, and (iv) existing methods for local error detection in related domains. We then outline where our sliding-window entropy framework fits within this broader literature, especially for mathematical OCR tasks.
1.1 Information Theory and OCR
Optical Character Recognition (OCR) can be analyzed through the prism of a noisy channel model. In such a view, an image $X$ (carrying text or mathematical symbols) is transmitted through a channel that might introduce noise via image blur, compression artifacts, or similar distortions. The output of that channel, a recognized text sequence $Y$, reflects an imperfect reconstruction of the original message. Foundational concepts[1] like entropy capture the overall uncertainty or “information content” in $X$ and $Y$. With higher resolution or lighter noise, $Y$ is more faithful to $X$ (i.e., $I(X;Y)$ is larger, and $H(Y\mid X)$ is smaller), implying fewer OCR errors.
Prior work has explored uncertainty quantification for text recognition using modern techniques like beam-search and dropout ensembles. For instance, Patel et al. [2] introduced Seq-UPS, a semi-supervised framework that leverages dropout-based uncertainty estimates to select reliable pseudo-labels, yielding improved OCR performance. Fang et al. [3] likewise proposed a Dropout Decoding method for large vision–language models; by explicitly discarding the most uncertain visual tokens, they reduced error rates on multimodal tasks such as visual question answering. Mathematical expressions, in particular, can be sensitive to OCR distortions: a single misread Greek symbol or missing bracket can disrupt an entire formula. Localizing such errors requires more fine-grained uncertainty signals, which motivates the present research.
1.2 GPT for Vision–Language Tasks
Generative pre-trained transformers (GPT) have evolved to handle multimodal inputs, merging deep text representations with image embeddings [4, 5]. Instead of restricting transformers to textual corpora, these extended architectures accept both an image $X$ and a textual prompt and generate an output token stream, $y_{1},y_{2},\dots,y_{n}$, from a vocabulary $\mathcal{V}$. Specifically,
$P(y_{i}\mid X,y_{1},\ldots,y_{i-1}),$
encodes the conditional probability of each new token given the image and preceding output tokens.
One promising application of these multimodal GPT models is mathematical OCR, where LaTeX transcription requires fine visual parsing. Luo and Xu [6] addressed contextual bias in transformer-based OCR decoders, emphasizing the importance of visual inputs for ambiguous symbols. Chen et al. [7] proposed a Recognition and Language Fusion Network that uses a lightweight language component alongside a vision core, significantly boosting math transcription accuracy on CROHME benchmarks. Similarly, Hyeon et al. [8] combined audio-based recognition with a small LM to rival GPT-4’s math OCR accuracy for lecture transcriptions. Going beyond purely mathematical text, Kim et al. [9] showed that large multimodal LLMs (e.g. GPT-4 Vision) can sometimes outperform classical OCR engines on difficult handwriting tasks, though integrative approaches that fuse external OCR signals can yield additional gains.
1.3 Token–Level Probability and Entropy
Modern large language model APIs commonly return a truncated set of top-$k$ tokens with associated log-probabilities at each time step. Converting this partial distribution into Shannon entropy (plus a single “tail” bucket for all lower-ranked tokens) offers a practical lower-bound estimate of the model’s true uncertainty at each position. This truncated entropy often suffices to detect peaks of uncertainty, especially if the tail probability is relatively small.
Recent advances in black-box LLM calibration suggest that model-generated confidences are themselves prone to miscalibration. Xiong et al. [10] found that naive confidence prompts in closed-source LLMs can be highly overconfident, while Yang et al. [11] proposed using multiple rephrasings of a query to improve answer consistency and estimate model uncertainty. Tonolini et al. [12] introduced Bayesian Prompt Ensembles to blend multiple prompts, producing more stable uncertainty measures with lower expected error. Yona et al. [13] further highlight that even aligned LLMs often sound more certain than their internal distribution suggests, while Wang et al. [14] provide a curated benchmark (UBench) to systematically test these models’ calibration on multiple-choice tasks.
Likewise for OCR, confidence gleaned from token-level distribution can guide error detection and correction. Although OCR often aggregated such signals over entire sequences, the ability to track local spikes in entropy opens clearer paths to targeted proofreading, as a single confusion can prove fatal for some mathematical expressions.
1.4 Localising Errors via Sliding Windows
In related fields like neural machine translation and speech transcription, researchers have experimented with short segment-level confidence measures to detect local mistakes. They often slice text into small windows and flag segments where the model’s probability mass is notably scattered. While this concept is relatively unexplored for GPT-based mathematical OCR, some studies demonstrate that local error-detection can substantially improve text quality across languages and scripts. For instance, Rigaud et al. [15] organized a post-OCR correction that split detection and correction into separate tasks, and Chen and Ströbel [16] proposed a hybrid method for post-processing outputs with a character-aware BERT (Bidirectional Encoder Representations from Transformers), pinpointing localized errors in historical handwriting. Iterative refinement [17] and two-stage pipelines [18] likewise demonstrate that targeting suspicious segments can outperform naive single-pass systems. Soper et al. [19] and Boros et al. [20] both show how large language models can be adapted for post-OCR text correction—though they also caution that out-of-the-box LLMs, while powerful, are not always reliable fixers of OCR errors. Bourne [21] made strides by synthesizing artificial errors to fine-tune a language model, cutting error rates even more effectively than some real-data approaches.
Within mathematical domains, local segmentation is especially relevant: a single bracket or minus sign can be misread, yet compromise the entire expression’s correctness. Our work introduces an entropy heat-mapping pipeline that computes a sliding-window average of per-token entropy. Regions with high mean entropy become hotspots for review—mirroring localized error detection in machine translation or historical text correction, but now tailored to GPT-based math OCR.
2 Problem Formulation
In this section we define the core task of detecting localised model uncertainty in GPT–based OCR outputs. Building on the broader information–theoretic view of OCR and conditional entropy, we now focus on per–token probability distributions and the goal of extracting short contiguous spans that exhibit exceptional uncertainty.
2.1 GPT–Based Mathematical OCR
We assume a vision–language model (e.g. GPT–4o) that ingests an input image $X$—potentially containing mathematical expressions—and decodes a token sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{n})$. Each token $y_{i}$ is chosen from a large vocabulary $\mathcal{V}$ according to the conditional distribution $P_{i}\bigl{(}y_{i}\mid X,y_{1},\dots,y_{i-1}\bigr{)}$. Production APIs expose log–probabilities for the $k$ most likely tokens, plus a tail bucket that aggregates the remaining probability mass.
At every position $i\in\{1,\dots,n\}$ we interpret the decoding step as a discrete random variable with entropy
$H_{i}\;=\;-\sum_{y\in\mathcal{V}}P_{i}(y)\,\log_{2}P_{i}(y),$ (1)
where $P_{i}(y)$ abbreviates $P_{i}\bigl{(}y\mid X,y_{1},\dots,y_{i-1}\bigr{)}$. Elevated values of $H_{i}$ signal that the model is unsure which token belongs at position $i$—a situation frequently arising when visual features are ambiguous (e.g. blurred superscripts) or when lexical context is rare.
2.2 Token–Level vs. Window–Level Entropy
Summing (1) over all positions yields a global measure of uncertainty, but a single scalar cannot reveal where errors are likely. Instead, we compute a sliding–window average. For a chosen window length $W$,
$A_{i}\;=\;\frac{1}{W}\;\sum_{j=i}^{i+W-1}H_{j},\qquad i=1,\dots,n-W+1.$ (2)
Spans whose mean entropy $A_{i}$ exceeds a rank– or threshold–based cutoff are flagged as entropy hotspots.
2.3 Hotspot Detection Objective
Given (2), the objective is to output a small set of candidate windows $\{[i_{1},i_{1}+W-1],\,[i_{2},i_{2}+W-1],\dots\}$ covering in total fewer than $M\!\cdot\!W$ tokens while capturing a disproportionately high fraction of the actual transcription errors. The intent is to direct human reviewers to those portions of the text that are most likely to contain mistakes, thereby reducing manual proofreading time.
2.4 Contrast with Global Entropy Approaches
Earlier work reported a single conditional–entropy value per page, facilitating comparisons across image resolutions but offering no localisation of errors. Our formulation provides:
  • a per–token truncated entropy $H_{i}$ (instead of a document–level aggregate);
  • a sliding–window average $A_{i}$ revealing short, highly uncertain segments; and
  • a procedure to extract the top–ranked hotspots for targeted review.
This fine granularity is crucial for mathematical OCR, where one errant symbol can invalidate an entire equation.
2.5 Key Model Assumptions
  1. 1.
    Top-$k$ Access. APIs provide only the highest-probability tokens; we approximate $H_{i}$ with the truncated‐entropy lower bound, sufficiently accurate when the tail mass is small.
  2. 2.
    Fixed Window Length $W$. A user-chosen $W$ trades off localisation precision against noise reduction; empirical tuning is discussed in Section 4.
  3. 3.
    Post-hoc Detection. Hotspots are identified after the entire sequence $\mathbf{y}$ has been generated, leaving streaming variants for future work.
With the problem precisely stated, Section 3 details the algorithm for computing window-level entropies, selecting the most uncertain spans, and optionally re-prompting the model to correct them.
3 Methodology
Building on the information–theoretic perspective and the formal task definition in Section 2, we now introduce our entropy heat–mapping pipeline. This section provides technical details on each step, including how we handle truncated probability distributions, define hotspots, and optionally re–prompt the model for error reduction.
3.1 Per–Token Entropy Extraction
Most commercial LLM endpoints, including GPT–4o, return the top–$k$ tokens and their log–probabilities at each decoding position. To compute the Shannon entropy at position $i$, we adopt the truncated entropy approximation described in Section 3.6.
Concretely, let
$\widehat{H}(i)\;=\;-\sum_{j=1}^{k}p^{(i)}_{j}\,\log_{2}p^{(i)}_{j}\;-\;p^{(i)}_{\mathrm{tail}}\,\log_{2}p^{(i)}_{\mathrm{tail}},$ (3)
where $p^{(i)}_{j}$ are the probabilities of the $k$ returned tokens and $p^{(i)}_{\mathrm{tail}}=1-\sum_{j=1}^{k}p^{(i)}_{j}$ collects the remaining mass.
Specifically:
  1. 1.
    We convert each returned log–probability to a probability $p^{(i)}_{j}=\exp(\log p^{(i)}_{j})$ for $j=1,\dots,k$.
  2. 2.
    We define the tail mass $p^{(i)}_{\mathrm{tail}}=1-\sum_{j=1}^{k}p^{(i)}_{j}$.
  3. 3.
    The truncated entropy $\widehat{H}(i)$ is obtained via Equation (3), which combines the $k$ explicit tokens plus one bucket representing all lesser probable tokens.
This yields a sequence $\{\widehat{H}(1),\ldots,\widehat{H}(n)\}$ of lower–bound entropies across the entire output. Although it underestimates the true entropy if the tail mass is sizable, we found in practice that $k=5$ or $k=10$ often suffices to capture the bulk of the distribution, making the gap negligible (we discuss implementation details in Section 3.4).
3.2 Sliding–Window Aggregation
Once $\widehat{H}(i)$ is available for $i\in\{1,\dots,n\}$, the goal is to highlight short contiguous sub–sequences that exhibit abnormally high average entropy. We employ the sliding–window average method, summarised here:
Window size $W$.
Given a fixed integer $W<n$, consider all windows $\,[i,i+W-1]\,$ of length $W$. Let
$S_{i}\;=\;\sum_{r=i}^{i+W-1}\widehat{H}(r),\quad A_{i}\;=\;\frac{S_{i}}{W},$ (4)
for $i=1,\dots,n-W+1$. We compute $A_{i}$ in one pass using the constant–time recurrence:
$S_{1}$ $=\sum_{r=1}^{W}\widehat{H}(r),\quad S_{i}=S_{i-1}+\widehat{H}(i+W-1)-\widehat{H}(i-1),$
$A_{i}$ $=\frac{S_{i}}{W}.$ (5)
The result is a signal $\{A_{1},\dots,A_{n-W+1}\}$ of length $n-W+1$.
Interpretation.
An elevated $A_{i}$ implies that positions $i$ through $i+W-1$ contain multiple tokens whose entropies spike simultaneously. When $W\!=\!1$, $A_{i}$ reduces to the single–token entropy $\widehat{H}(i)$, capturing the most granular level of pointwise uncertainty. Larger $W$ identifies sub–sequences of moderate length where the model is consistently uncertain, possibly reflecting more complex OCR challenges like multiline equations or complicated notation contexts.
3.3 Hotspot Definition and Visualisation
High–entropy windows (“hotspots”) are those intervals where $A_{i}$ exceeds some threshold. We adopt two broad strategies:
  • Rank–Based Selection. Sort the windows by descending $A_{i}$ and pick the top $M$, for a chosen $M$ (e.g. $M=5$). This ensures that only a handful of segments are flagged regardless of the absolute scale of entropies.
  • Percentile Thresholds. Calculate the $\alpha$–th percentile (e.g. $\alpha=90$) of $\{A_{1},\dots,A_{n-W+1}\}$, and mark any window whose $A_{i}$ is above that cutoff. This dynamic approach adapts to different distributions of entropies on each document.
We then generate a heatmap or color-coded overlay on the final transcribed text, where each token is shaded proportional to its local mean entropy. Consecutive tokens with high shading typically cluster into a single uncertain region; these clusters can be examined by an OCR reviewer for potential errors such as missing parentheses or garbled symbols.
3.4 Implementation Details
Our prototype is written in Python and relies on the GPT–4o API. Key parameters:
[…]
  • Entropy Computation. We accumulate truncated entropy via Eq. (3) for each token. Special tokens (e.g. line–breaks, code fence markers) can be excluded if desired, but in our proof–of–concept we typically keep them so that missing punctuation might also be revealed.
  • Windowing. We select $W\in\{5,10,20\}$ in different runs, verifying quickly that $W=10$ is an acceptable compromise between making hotspots stable (not just single unlucky tokens) and still local enough for error detection.
  • Complexity. The entire pipeline runs in $O(n)$ time once the probabilities are retrieved, so even multi–page documents (thousands of tokens) are feasible. Our proof–of–concept dataset remains small, but the method should scale to large volumes of scientific text.
3.4.1 Algorithm
Input: scanned page path $image\_path$;
top–$k$ alternatives $k$ (log-probabilities);
sliding-window length $W$;
number of windows to report $M$;
VLM identifier (gpt_4o).
Output: LaTeX transcript $T$;
per-token truncated entropies $\widehat{H}(1{:}n)$;
the $M$ windows with highest mean entropy.
1
21. Pre–process;
3 Parse CLI/GUI arguments $(image\_path,k,W,M)$;
4 if file_missing$(image\_path)$ then exit;
5  // Abort if missing Encode image in Base64 for API transmission;
6
72. Prompt and inference;
8 Build system and user prompts (image + instructions);
9 Call GPT–4o with logprobs=True, top_logprobs=$k$;
10 Receive best transcript $T$ and token log-probs $\ell_{i,j}$;
11
123. Token-level truncated entropy;
13 for $i\leftarrow 1$ to $n$ do // $i=1{:}n$ tokens
14       $p_{i,j}\leftarrow e^{\ell_{i,j}}\quad(j=1{:}k)$;
15       $p_{\text{tail}}\leftarrow 1-\sum_{j=1}^{k}p_{i,j}$;
16       $\widehat{H}(i)\leftarrow-\sum_{j=1}^{k}p_{i,j}\log_{2}p_{i,j}-p_{\text{tail}}\log_{2}p_{\text{tail}}$;
17
18
194. Sliding-window aggregation;
20 $S\leftarrow\sum_{r=1}^{W}\widehat{H}(r)$;
21 for $i\leftarrow 1$ to $n-W+1$ do
22       $A_{i}\leftarrow S/W$;
23       $S\leftarrow S+\widehat{H}(i+W)-\widehat{H}(i)$;
24
25
265. Hotspot selection;
27 Use a max-heap to return the $M$ indices with largest $A_{i}$;
28
296. Output;
30 Save full LaTeX document $T$ next to $image\_path$;
31 Return $\bigl{(}T,\widehat{H}(1{:}n),\{A_{i}\},\text{top-}M\text{ windows}\bigr{)}$;
32
337. (Optional) post-processing;
34 Highlight hotspots in $T$ or re-prompt GPT on each suspicious window;
Algorithm 1 GPT–based OCR with Sliding–Window Entropy Hotspots
3.5 Optional Re–Prompt or Post–Editing Loop
Finally, although our primary goal is to locate potential errors, one can incorporate a correction stage:
  1. 1.
    Identify windows with highest $A_{i}$.
  2. 2.
    Extract the associated text snippet as a query.
  3. 3.
    Prompt GPT–4o (or a specialized LaTeX editing system) to re–examine that snippet in context and propose corrections.
Alternatively, a human operator can revise only these suspicious regions, thus reducing effort. While the experiments in Section 4 focus on human annotation, such automated re–prompting is a natural extension for large–scale digitisation pipelines where fully manual review would be time–consuming.
3.6 Top-$k$ + Tail Entropy is a Lower Bound
In our framework, we approximate the full token distribution at each position by tracking the top $k$ tokens plus a single “tail” bucket that merges the probabilities of all remaining tokens. Intuitively, ignoring the distinctions among tail tokens can only reduce the entropy value—leading to a lower bound on the true Shannon entropy. Below, we provide a self-contained proof of this fact.
3.6.1 Problem Setting
Let $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$ be the model’s full vocabulary of $n$ tokens, and let $X$ be the random variable denoting which token is actually emitted at a given decoding position. We write
$p_{i}\;=\;\Pr[X=v_{i}],\quad p_{i}\geq 0,\quad\sum_{i=1}^{n}p_{i}=1\,.$
Suppose we know only the probabilities of the $k$ most likely tokens (without loss of generality $v_{1},\dots,v_{k}$), and define their probabilities in descending order:
$p_{1}\,\geq\,p_{2}\,\geq\,\dots\,\geq\,p_{k},\quad\text{with }k<n.$
Let
[…]
The true Shannon entropy of $X$ is
$H_{\mathrm{true}}(X)\;=\;-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}.$
3.6.2 Approximate Top-$k$ + Tail Entropy
We form a new random variable $Y$ that retains the identity of the top-$k$ tokens and merges the others into one “tail” symbol:
$Y\;=\;\begin{cases}v_{i},&\text{if }X=v_{i}\text{ for }i\in\{1,\dots,k\},\\ \text{\it tail},&\text{otherwise}.\end{cases}$
Hence,
$\Pr[Y=v_{i}]=p_{i}\quad(i=1,\dots,k),\quad\text{and}\quad\Pr[Y=\text{\it tail}]=p_{\mathrm{tail}}.$
We then compute the entropy of $Y$:
$\widehat{H}\;=\;H(Y)\;=\;-\biggl{(}\sum_{i=1}^{k}p_{i}\log_{2}p_{i}\;+\;p_{\mathrm{tail}}\log_{2}p_{\mathrm{tail}}\biggr{)}.$
Clearly, $\widehat{H}$ is simpler to calculate than the true entropy $H_{\mathrm{true}}(X)$ because it treats the entire set of lower-probability tokens as a single event. We now verify that this approximation cannot exceed the actual entropy of $X$.
3.6.3 Claim and Proof
Claim:
$\widehat{H}\;\;=\;\;H(Y)\;\;\leq\;\;H_{\mathrm{true}}(X)\;\;=\;\;-\sum_{i=1}^{n}p_{i}\,\log_{2}p_{i}.$
Equality holds if and only if $p_{\mathrm{tail}}=0$ (i.e. the top $k$ tokens already sum to probability 1).
Proof Sketch.
Let $f$ be the deterministic mapping $X\mapsto Y$, which sends $v_{i}$ to itself for $i\leq k$ and all other tokens $v_{k+1},\dots,v_{n}$ to the single symbol “tail.” From standard chain-rule properties of entropy, we have
$H(X,Y)\;=\;H(X)\,=\;H(Y)+H(X\mid Y).$
Since $Y=f(X)$ is a function of $X$, we know $H(X\mid Y)\geq 0$ with equality if and only if there is a one-to-one correspondence between $X$ and $Y$ [1]. Combining this with (1) yields
[…]
Thus, $\widehat{H}=H(Y)$ never exceeds the true entropy. If $p_{\mathrm{tail}}>0$, then at least two distinct tokens map to tail, implying $H(X\mid Y)>0$ and a strict inequality. Conversely, if $p_{\mathrm{tail}}=0$, all probability mass lies in the top $k$ tokens and $f$ is effectively the identity map, giving $H(X\mid Y)=0$ and $\widehat{H}=H_{\mathrm{true}}(X)$.
This elementary argument justifies our assertion in Section 3.1 that the top-$k$+tail entropy approximation cannot overshoot the real entropy.
In the next section, we detail our proof–of–concept corpus and annotation procedure to evaluate whether these entropy–based hotspots indeed coincide with model transcription errors in mathematical OCR tasks.
4 Experimental Setup (Proof-of-Concept)
We now describe our small-scale experiment designed to test whether sliding-window entropy hotspots effectively highlight local OCR errors. In contrast to large benchmarking studies, we focus on a proof-of-concept that relies on human inspection and qualitative judgments to see if high-entropy spans correlate with actual mistakes in GPT-based mathematical OCR.
4.1 Mini-Corpus
We curated a modest collection of 12 scanned pages from arXiv preprints that contain a variety of mathematical expressions, ranging from inline equations (single-line formulas with moderate symbol complexity) to multi-line displays (with fractions, sums, and Greek letters). Each page was rendered to an image at three distinct resolutions: 72 dpi, 150 dpi, and 300 dpi. This yields a total of 36 image instances.
On these pages, the text length varies between 200–600 tokens, with mathematical notations interspersed among paragraphs. Although a more comprehensive survey could include dozens or hundreds of pages, our goal here is simply to demonstrate feasibility and gather initial human feedback on the proposed approach.
4.2 Annotation Protocol
In implementing the pipeline for this experiment, we followed the pseudo-code in Algorithm 1 (shown earlier) to produce both the OCR transcript and its corresponding token-level entropies. Concretely, we (1) invoked a GPT-4o endpoint with image-containing prompts to obtain token-by-token log-probabilities; (2) computed truncated entropy at each token position; then (3) generated the sliding-window averages that identified high-entropy “hotspots” for review. Specifically:
  1. 1.
    Transcript Generation. For each of the 36 images, we used GPT–4o with a standard prompt requesting faithful LaTeX transcription of any recognized formulas. We stored the complete token-by-token output and its top-$k$ log-probability distributions.
  2. 2.
    Entropy Computation. Following Section 3.1 and the pseudo-code steps (Algorithm 1, Steps 4–5), we computed truncated entropies $\widehat{H}(i)$ for each token $i$. Then we generated sliding-window averages $A_{i}$ in windows of length $W=10$.
  3. 3.
    Hotspot Identification. We sorted all windows by descending $A_{i}$ and selected the top 3 windows as hotspots. Each hotspot is thus a contiguous 10-token segment suspected to contain OCR errors.
  4. 4.
    Manual Error Labeling. Three expert annotators (graduate students in computer science) compared each transcript to the displayed math on the original PDF. They flagged any text token that omitted, replaced, or garbled a mathematical symbol (e.g., a subscript turned into a digit) or revealed major syntax breaks (e.g., unmatched braces). The same marking guidelines were applied across all resolution variants.
  5. 5.
    Hotspot–Error Overlap. Finally, a simple script compared the annotator-flagged tokens with the entropy hotspots for each transcript, recording whether the tokens fell inside or outside the top-3 windows.
Since our goal was not to compile precise metrics but rather to gain qualitative insight, we did not require strict inter-annotator agreement. Instead, each individual’s markings were tallied separately, and any token flagged as erroneous by any of the three annotators was considered a failure in the transcript. In the next section (Section 5), we outline the principal outcomes, including sample heatmaps and anecdotal observations on each resolution’s performance.
4.3 Evaluation Procedure
To keep the study proof-of-concept, we avoided complicated quantitative measures. Instead, the procedure emphasized ease of review and interpretability:
  • Side-by-Side Interface. Annotators saw two panels: (1) the original scanned image with math formulas, and (2) the GPT–4o LaTeX output, where tokens belonging to each hotspot were visually highlighted in color.
  • Qualitative Feedback. The reviewers provided free-text comments on whether the hotspots did or did not capture real misread tokens. They could also note any false alarms (windows with high entropy but no actual error).
  • Time Estimates. Each annotator reported approximate time spent (±$(\pm 15\,\text{seconds})$ on each transcript. Although these reports are subjective, they confirm whether focusing on hotspots reduced their proofing effort in practice.
In choosing only the top 3 windows per page, we deliberately limited the fraction of text to be flagged, thus simulating a real-world need to minimize human review. Our primary question was whether these few hotspots would capture the majority of recognized errors. The next section (Section 5) gives the principal outcomes, including sample heatmaps and anecdotal observations on each resolution’s performance.
5 Findings
In this section, we describe the main outcomes of our proof-of-concept study on localised entropy hotspots in GPT-based OCR. We first provide qualitative evidence that visually identifiable hotspots do in fact coincide with real transcription errors. Next, we highlight specific types of mistakes uncovered and consider annotator feedback on how the hotspots guide human correction. Finally, we discuss preliminary observations about window size and resolution variations.
5.1 Visual Alignment Between Hotspots and Errors
Despite having no direct knowledge of the ground-truth tokens, the entropy-based approach (Section 3) consistently flagged local segments in which GPT–4o struggled, often suggestive of a morphological or typographical confusion. When displaying the text with shaded tokens (darkest shades corresponding to highest average entropy), our three annotators noted that actual mistakes—such as a Greek letter incorrectly recognized as an English letter—tended to cluster within these dark regions.
For instance, on pages containing multiline equations, three or more tokens near the fraction bar were systematically marked as high entropy, likely reflecting GPT–4o’s uncertainty in parsing subtle visual cues around fraction boundaries or in discerning subscript positions. In the annotated transcripts, nearly all misread symbols within these equations fell squarely under these high-entropy windows. Although some entropy peaks contained no “catastrophic” error, our annotators observed that the text within such windows still lacked clarity or consistency (e.g. spurious brace insertion).
5.2 Case Studies: Types of Errors Found
Across the 36 annotated transcripts (12 pages $\times$ 3 resolutions), we can group salient OCR mistakes into four representative categories:
  1. 1.
    Minus–Sign Ambiguities. A minus sign rendered in a low-resolution image sometimes resembled a tilde or dash, producing tokens that GPT–4o assigned moderate probabilities to a variety of possible characters. The resulting hotspot regions often contained one or two incorrect symbols that changed the equation.
  2. 2.
    Mismatched Delimiters. Parentheses or braces, when partially hidden or incorrectly aligned, triggered high local entropy. In multiple examples, the text around an unmatched opening brace also included partially garbled subsequent tokens.
  3. 3.
    Greek Letter Confusions. Letters like $\pi$ and $\rho$ can be mistaken for English ‘p‘ or ‘n‘. Such confusion led to upward spikes in entropy whenever the contextual cues about whether the symbol was a math variable or plain text were weak (e.g. a small snippet of math inline with normal prose).
  4. 4.
    Symbol Dropping. Entire markers such as underscores (for subscripts) sometimes vanished at lower resolutions, again strongly increasing GPT–4o’s local uncertainty. The missing symbol also tended to shift subsequent tokens, compounding the script confusion.
Though not exhaustive, these examples underscore that the sliding–window method successfully funnels attention toward typical failure modes of math OCR. In a more automated pipeline, it could function as a triage mechanism for challenging pages.
5.3 Reviewer Effort
Our human annotators reported that spotting error-prone tokens became faster when the three highest-entropy windows were highlighted in color. Annotators repeatedly skipped those portions of the text with low entropy, focusing their scrutiny on only 10–15 % of the tokens. While we did not attempt formal time–tracking, multiple participants commented that having a “quick visual highlight” prevented them from manually rereading entire pages for minor mistakes. Therefore, even without advanced metrics, the hotspots appear to reduce the overhead of verifying GPT’s output.
The main frustration noted was that some windows were “false alarms,” in the sense that no clear error was discovered there. These were often short sequences of punctuation near math expressions; the model seemingly was uncertain but did not produce an actual error. Targeted re-prompting (Section 3.5) might resolve these borderline cases.
5.4 Observations on Window Size and Resolution
Finally, we examined how the window length $W$ and the input resolution interact in producing hotspots:
Window Length.
In a pilot trial with $W=1$, numerous isolated tokens scored high entropy, but the approach sometimes fragmented a single error across several micro–windows. At the other extreme, a large window ($W=20$) could blur distinct issues together (e.g. a missing brace in one equation and a Greek letter error in another). Our final choice, $W=10$, appeared to strike a suitable balance, grouping related tokens without spanning too many lines of text.
Resolution.
Unsurprisingly, the 72 dpi transcripts included more frequent hotspots, reflecting the model’s overall difficulty in interpreting blurry details. Meanwhile, for images at 300 dpi, the number of high-entropy windows typically dropped by 80% or more.
In short, the findings confirm that our sliding-window entropy heatmaps highlight local error-prone sections of GPT-based OCR output. Although some caution is warranted around false alarms and parameter choices like $W$, this approach holds promise for faster, more focused review of machine-transcribed mathematical content.
6 Discussion
Our proof-of-concept study shows that sliding-window entropy can direct human reviewers toward local regions of GPT–based OCR output where recognition errors are most likely to occur. In contrast to prior, global entropy approaches that yield a single scalar per page, our method yields a fine-grained “heat map” over the output tokens, identifying sub-sequences (hotspots) with heightened uncertainty.
6.1 Why Local Entropy Peaks Reflect Actual Error Zones
At a high level, an elevated local entropy arises when the model’s probability mass is spread over multiple competing tokens. Such competition frequently stems from ambiguous visual cues (e.g. partial subscript, unclear symbol boundaries) or from parsing difficulties (e.g. a suspicious brace or bracket). The data indicate that these same ambiguous visual features are precisely where practical OCR mistakes occur: misread minus signs, mismatched parentheses, or dropped subscripts. Consequently, windowed entropy acts as a simple proxy for the model’s internal hesitation—highlighting precisely those tokens that “look suspicious” to GPT–4o.
Furthermore, our qualitative review suggests that, even if the system does not always produce a devastating error at each high-entropy position, the suspicious region often contains subtle misalignments or unusual tokens that warrant a closer human look. This is especially relevant in mathematical documents, where even minor symbol confusion can derail an entire expression.
6.2 Integration into Digitisation Pipelines
In practical workflows (e.g. arXiv paper ingestion or scientific scanning in digital libraries), volume is large and manual proofreading resources are limited. Our results show that focusing reviewer effort on a handful of high-entropy windows can reduce the necessity of line-by-line scanning. Implemented as a postprocessing step, the proposed module could:
  1. 1.
    Read Output Logprobs: Extract truncated probabilities from the GPT model as it decodes each text token.
  2. 2.
    Compute Windowed Entropies: Obtain $A_{i}$ for each sub-sequence of length $W$ (Sections 3.23.3).
  3. 3.
    Highlight or Re-Prompt: Mark the top-$M$ windows in the final transcript or pass them back to GPT–4o for a localized re-check.
Notably, the entire process is straightforward to script, requiring $O(n)$ time once per-token log-probabilities are available. This overhead seems negligible compared to the time saved in human correction.
6.3 Limitations of Qualitative Conclusions
Because our experiments relied mostly on human inspection of a small sample of pages, we acknowledge that the present results do not yield rigorous quantitative benchmarks, such as recall or precision rates for spotting errors. Instead, we have offered a “first evidence” that local entropy spikes do align with the problematic tokens in practice. Further expansions of the dataset, or a more systematic labeling of errors at scale, could confirm these findings under broader conditions.
Additionally, we have not extensively dissected false-positive hotspots (entropy spikes that do not contain real errors). In the handful of such windows we observed, the model was apparently uncertain but still arrived at correct symbols. Tailoring the method to further reduce these spurious alerts might require additional heuristics or re-prompting strategies.
6.4 Broader Implications
Despite these limitations, our findings resonate with the broader trend of integrating fine-grained uncertainty measures into large language model pipelines. By focusing on local pockets of doubt, end users can apply targeted verification or domain-specific knowledge, boosting both efficiency and trustworthiness. We see parallels in other multimodal tasks: for instance, scanning a neural speech recognizer’s token-level entropies can likewise flag unclear utterances for a second pass. Extending such local entropy maps to more advanced tasks—like multi-column page layouts, embedded figures, or even interactive editing—presents exciting directions for future work.
In conclusion, these observations reinforce the general principle that “where the model is least certain” often overlaps with “where real OCR errors lurk.” The sliding-window approach, while simple, appears surprisingly effective as a lightweight mechanism to triage GPT-based OCR outputs for potential mistakes.

### v2

@@
 Abstract
-Vision–language models such as OpenAI GPT-4o can transcribe mathematical documents directly from images, yet their token-level confidence signals are seldom used to pinpoint local recognition mistakes.
-We present an entropy-heat-mapping proof-of-concept that turns per-token Shannon entropy into a visual “uncertainty landscape.” By scanning the entropy sequence with a fixed-length sliding window, we obtain hotspots that are likely to contain OCR errors such as missing symbols, mismatched braces, or garbled prose.
+Vision-language models such as OpenAI GPT-4o can transcribe mathematical documents directly from images, yet their token-level confidence signals are seldom used to pinpoint local recognition mistakes.
+We present an entropy-heat-mapping proof-of-concept that turns per-token Shannon entropy into a visual ”uncertainty landscape”.
+By scanning the entropy sequence with a fixed-length sliding window, we obtain hotspots that are likely to contain OCR errors such as missing symbols, mismatched braces, or garbled prose.
 Using a small, curated set of scanned research pages rendered at several resolutions, we compare the highlighted hotspots with the actual transcription errors produced by GPT-4o.
 Our analysis shows that the vast majority of true errors are indeed concentrated inside the high-entropy regions.
-This study demonstrates—in a minimally engineered setting—that sliding-window entropy can serve as a practical, lightweight aid for post-editing GPT-based OCR.
-All code, sample data, and annotation guidelines are released to encourage replication and further research.
-1 Background and Related Work
-This section provides context on (i) the information–theoretic lens for understanding OCR, (ii) the extension of GPT from purely text-based to vision–language modeling, (iii) prior efforts leveraging per-token probability distributions for confidence checks, and (iv) existing methods for local error detection in related domains.
-We then outline where our sliding-window entropy framework fits within this broader literature, especially for mathematical OCR tasks.
+This study demonstrates–in a minimally engineered setting–that sliding-window entropy can serve as a practical, lightweight aid for post-editing GPT-based OCR.
+All code and annotation guidelines are released to encourage replication and further research.
+Keywords: Sliding-window entropy, Mathematical OCR, Token-level uncertainty heat-mapping, Error localisation.
+1 Introduction and Related Work
+Transcribing mathematical documents from scanned images poses unique challenges, as subtle symbol confusion (e.g., minus signs vs.
+tildes) can undermine entire expressions.
+In this section, we review three broad research directions: (i) information–theoretic perspectives on OCR, (ii) the emergence of GPT for multimodal tasks, and (iii) prior efforts at uncertainty-based error detection—both global and local.
+We then show how our sliding-window entropy framework extends these lines of work to pinpoint where within a GPT-based transcript recognition mistakes arise, with a particular focus on mathematical OCR.
 1.1 Information Theory and OCR
-Optical Character Recognition (OCR) can be analyzed through the prism of a noisy channel model.
-In such a view, an image $X$ (carrying text or mathematical symbols) is transmitted through a channel that might introduce noise via image blur, compression artifacts, or similar distortions.
-The output of that channel, a recognized text sequence $Y$, reflects an imperfect reconstruction of the original message.
-Foundational concepts[1] like entropy capture the overall uncertainty or “information content” in $X$ and $Y$.
-With higher resolution or lighter noise, $Y$ is more faithful to $X$ (i.e., $I(X;Y)$ is larger, and $H(Y\mid X)$ is smaller), implying fewer OCR errors.
-Prior work has explored uncertainty quantification for text recognition using modern techniques like beam-search and dropout ensembles.
+Optical Character Recognition (OCR) can be framed through the lens of a noisy-channel model [1], where an image $X$ containing text must be decoded despite potential blur, compression artifacts, or other noise sources—resulting in an imperfect output sequence $Y$.
+Classical information measures (e.g., entropies $H(X)$ and $H(Y)$, or the mutual information $I(X;Y)$) help describe how uncertainty in $X$ propagates to $Y$.
+Modern research has further explored ways to quantify OCR uncertainty and mitigate errors.
 For instance, Patel et al.
-[2] introduced Seq-UPS, a semi-supervised framework that leverages dropout-based uncertainty estimates to select reliable pseudo-labels, yielding improved OCR performance.
+[2] derive per-sequence uncertainty estimates via Monte Carlo dropout to filter unreliable pseudo-labels in text recognition.
 Fang et al.
-[3] likewise proposed a Dropout Decoding method for large vision–language models; by explicitly discarding the most uncertain visual tokens, they reduced error rates on multimodal tasks such as visual question answering.
-Mathematical expressions, in particular, can be sensitive to OCR distortions: a single misread Greek symbol or missing bracket can disrupt an entire formula.
-Localizing such errors requires more fine-grained uncertainty signals, which motivates the present research.
+[3] describe a dropout-decoding strategy in large vision–language models, illustrating that high global uncertainty frequently correlates with poor recognition accuracy.
+These approaches demonstrate that confidence measures can be leveraged to improve OCR outputs—particularly for complex tasks like mathematical symbol detection, where even a single misread bracket may invalidate an entire formula.
 1.2 GPT for Vision–Language Tasks
-Generative pre-trained transformers (GPT) have evolved to handle multimodal inputs, merging deep text representations with image embeddings [4, 5].
-Instead of restricting transformers to textual corpora, these extended architectures accept both an image $X$ and a textual prompt and generate an output token stream, $y_{1},y_{2},\dots,y_{n}$, from a vocabulary $\mathcal{V}$.
-Specifically,
-$P(y_{i}\mid X,y_{1},\ldots,y_{i-1}),$
-encodes the conditional probability of each new token given the image and preceding output tokens.
-One promising application of these multimodal GPT models is mathematical OCR, where LaTeX transcription requires fine visual parsing.
-Luo and Xu [6] addressed contextual bias in transformer-based OCR decoders, emphasizing the importance of visual inputs for ambiguous symbols.
-Chen et al.
-[7] proposed a Recognition and Language Fusion Network that uses a lightweight language component alongside a vision core, significantly boosting math transcription accuracy on CROHME benchmarks.
-Similarly, Hyeon et al.
-[8] combined audio-based recognition with a small LM to rival GPT-4’s math OCR accuracy for lecture transcriptions.
-Going beyond purely mathematical text, Kim et al.
-[9] showed that large multimodal LLMs (e.g.
-GPT-4 Vision) can sometimes outperform classical OCR engines on difficult handwriting tasks, though integrative approaches that fuse external OCR signals can yield additional gains.
-1.3 Token–Level Probability and Entropy
-Modern large language model APIs commonly return a truncated set of top-$k$ tokens with associated log-probabilities at each time step.
-Converting this partial distribution into Shannon entropy (plus a single “tail” bucket for all lower-ranked tokens) offers a practical lower-bound estimate of the model’s true uncertainty at each position.
-This truncated entropy often suffices to detect peaks of uncertainty, especially if the tail probability is relatively small.
-Recent advances in black-box LLM calibration suggest that model-generated confidences are themselves prone to miscalibration.
-Xiong et al.
-[10] found that naive confidence prompts in closed-source LLMs can be highly overconfident, while Yang et al.
-[11] proposed using multiple rephrasings of a query to improve answer consistency and estimate model uncertainty.
-Tonolini et al.
-[12] introduced Bayesian Prompt Ensembles to blend multiple prompts, producing more stable uncertainty measures with lower expected error.
-Yona et al.
-[13] further highlight that even aligned LLMs often sound more certain than their internal distribution suggests, while Wang et al.
-[14] provide a curated benchmark (UBench) to systematically test these models’ calibration on multiple-choice tasks.
-Likewise for OCR, confidence gleaned from token-level distribution can guide error detection and correction.
-Although OCR often aggregated such signals over entire sequences, the ability to track local spikes in entropy opens clearer paths to targeted proofreading, as a single confusion can prove fatal for some mathematical expressions.
-1.4 Localising Errors via Sliding Windows
-In related fields like neural machine translation and speech transcription, researchers have experimented with short segment-level confidence measures to detect local mistakes.
-They often slice text into small windows and flag segments where the model’s probability mass is notably scattered.
-While this concept is relatively unexplored for GPT-based mathematical OCR, some studies demonstrate that local error-detection can substantially improve text quality across languages and scripts.
-For instance, Rigaud et al.
-[15] organized a post-OCR correction that split detection and correction into separate tasks, and Chen and Ströbel [16] proposed a hybrid method for post-processing outputs with a character-aware BERT (Bidirectional Encoder Representations from Transformers), pinpointing localized errors in historical handwriting.
-Iterative refinement [17] and two-stage pipelines [18] likewise demonstrate that targeting suspicious segments can outperform naive single-pass systems.
-Soper et al.
-[19] and Boros et al.
-[20] both show how large language models can be adapted for post-OCR text correction—though they also caution that out-of-the-box LLMs, while powerful, are not always reliable fixers of OCR errors.
-Bourne [21] made strides by synthesizing artificial errors to fine-tune a language model, cutting error rates even more effectively than some real-data approaches.
-Within mathematical domains, local segmentation is especially relevant: a single bracket or minus sign can be misread, yet compromise the entire expression’s correctness.
-Our work introduces an entropy heat-mapping pipeline that computes a sliding-window average of per-token entropy.
-Regions with high mean entropy become hotspots for review—mirroring localized error detection in machine translation or historical text correction, but now tailored to GPT-based math OCR.
+Over the past few years, generative pre-trained transformers have evolved from purely text-based systems into multimodal networks that interpret images alongside textual prompts [4, 5].
+These extended GPT engines map image embeddings and partially decoded text into probabilities over a large vocabulary.
+Vision–language GPT models have proven especially promising for mathematical OCR, where long-range context is needed to disambiguate complex notation.
+Luo and Xu [6], for example, emphasize the role of precise visual cues to differentiate visually similar Greek letters, while Chen et al.
+[7] integrate a lightweight language model with a dedicated visual core to reduce transcription errors on CROHME math benchmarks.
+Recent studies also show that GPT’s visual modules can handle challenging inputs like handwritten math [8] or audio-based lecture notes [9], sometimes outperforming traditional OCR pipelines.
+Nevertheless, even top-tier vision–language models can generate localized mistakes—missing or confusing symbols—in ways that standard system-level metrics fail to capture.
+1.3 Token-Level Entropy versus Global Uncertainty
+When large language models (LLMs) generate text, their APIs often provide the top-$k$ tokens with corresponding log-probabilities at each decoding step.
+Summing across these probabilities yields a truncated distribution suitable for approximate Shannon entropy.
+Several works have noted that peaks in token-level entropy can reflect points of confusion in text generation [10, 11, 12, 13, 14].
+For OCR, such confusion likely stems from visually ambiguous inputs: a blurred superscript, a partially occluded minus sign, and so on.
+Yet, many existing systems focus on global or sequence-level confidence.
+For instance, Wang et al.
+[15] calibrate a single “energy-based” score for visual classifiers that flags uncertain images, while Chun et al.
+[16, 17] embed entire image–text pairs into probabilistic spaces and measure differential entropy as a reliability indicator.
+Similarly, Upadhyay et al.
+[18, 19] attach Bayesian adapters to frozen networks, using overall negative log-evidence or variance as a global confidence metric to detect untrustworthy predictions.
+Fei et al.
+[20] rank candidate captions by their total sequence-level entropy to identify robust image captions, and Ji et al.
+[21] monitor global KL-divergence to decide when multimodal embeddings are stable.
+For cross-modal retrieval tasks, Li et al.
+[22] reduce all Dirichlet-based evidence into a single score, and Khan & Fu [23] show that the overall distributional entropy correlates with unreliable VQA results.
+Zhou et al.
+[24] also rely on total entropy to spot ambiguous event-based actions.
+While these global measures can signal that an entire output is potentially flawed, they provide no direct insight on which part of the text is at fault.
+By contrast, we exploit local per-token signals to automatically pinpoint suspicious sub-sequences—crucial for post-editing mathematical text, where minor symbol mismatches can cause outsized damage.
+1.4 Local Error Detection via Sliding Windows
+To address the problem of where errors occur, researchers in machine translation or speech recognition often slice transcripts into short windows and measure local confidence.
+Rigaud et al.
+[25] propose a post-OCR correction process that operates on small segments, while Chen and Ströbel [26] refine OCR output by selectively re-checking suspicious tokens with a character-aware language model.
+Similar segment-wise or iterative approaches have been adopted for historical text [27, 28], and indeed large language models like BART have been retrained to patch local OCR mistakes [29, 30].
+Bourne [31] goes further by synthesizing artificial errors in order to robustly fine-tune correction models, reducing error rates more effectively.
+Notably, while these pipelines vary, they share a common theme: error-prone sub-sequences can be more effectively targeted—either for human proofing or re-prompted corrections—than scanning an entire document blindly.
+In the context of mathematical OCR, local segmentation is especially critical: dropping a subscript underscore or mixing up braces can invalidate a formula.
+Our work introduces entropy heat-mapping as a lightweight technique to identify hotspots of local uncertainty.
+Specifically, we (i) compute truncated entropy for each token, (ii) apply a sliding window to produce interval-based averages, and (iii) highlight the windows with highest mean entropy.
+These windows correlate strongly with actual symbol confusion or missing tokens, as demonstrated in our proof-of-concept results.
+1.5 Our Contribution
+Although several recent studies [15, 16, 18, 20, 21, 22, 23, 24] employ global entropy or other confidence scores to gauge the overall reliability of vision–language models, the question of where recognition errors arise inside a generated transcript remains largely unexplored.
+The present paper addresses that gap by proposing an entropy-based localisation framework for GPT-driven mathematical OCR.
+Our main contributions are:
+  1. 1.
+    Sliding-window Shannon analysis.
+We transform the per-token log-probabilities returned by modern GPT APIs into a linear-time, sliding-window entropy signal that yields a fine-grained “uncertainty landscape” over the output text.
+  2. 2.
+    Entropy heat-mapping for error detection.
+Visualising the windowed entropies highlights short contiguous spans—hotspots—whose elevated uncertainty correlates strongly with actual transcription mistakes such as missing symbols, mismatched braces, or dropped subscripts.
+  3. 3.
+    Instantiating the classical coarse-graining property for top-$k$+tail entropy.
+It has long been known in information theory (e.g.
+Cover & Thomas, Sec.
+2) that coarse-graining—merging several outcome categories into one—can only decrease Shannon entropy.
+We adapt this classical fact to the practical GPT setting in which APIs expose probabilities only for the top-$k$ tokens, collapsing the remaining mass into a single “tail” bucket.
+Our short, self-contained derivation shows exactly how the coarse-graining inequality applies to per-token log-probabilities and clarifies its implications for black-box uncertainty extraction in vision–language OCR.
+  4. 4.
+    Proof-of-concept human evaluation.
+On a curated mini-corpus of scanned research pages rendered at three resolutions (72, 150, and 300 dpi), we demonstrate that the vast majority of human-verified OCR errors fall inside the few highest-entropy windows, validating the practical utility of the proposed heat-mapping approach.
+  5. 5.
+    Open-source resources.
+All code, sample images, entropy-extraction scripts, and detailed annotation guidelines are publicly released to encourage replication and further research on local uncertainty cues in multimodal GPT systems.
+In essence, the paper converts raw probabilistic signals—already available from mainstream GPT endpoints—into an actionable tool for post-editing: it pinpoints precisely where the model is least certain and therefore where human reviewers or automatic re-prompting should focus.
+Reproducibility.
+The complete implementation and annotation protocol are freely available at https://github.com/Alexei-WLU/scan2latex-entropy.
+1.6 Organization of the Paper
+We formally define the local uncertainty detection problem in Section 2, specifying how GPT-based OCR outputs and per-token probabilities lead to window-level entropies.
+Next, Section 3 presents our sliding-window technique, including exact computation of top-$k$+tail entropy and approaches for highlighting hotspots in linear time.
+In Section 4, we detail the proof-of-concept experiments and human annotation protocol, while Section 5 reports qualitative and quantitative results.
+We discuss broader implications and limitations in Section 6, and Section 7 concludes with future prospects for scaling entropy heat-mapping to real-world OCR scenarios.
 2 Problem Formulation
@@
   1. 1.
-    Top-$k$ Access. APIs provide only the highest-probability tokens; we approximate $H_{i}$ with the truncated‐entropy lower bound, sufficiently accurate when the tail mass is small.
+    Top-$k$ Access. APIs provide only the highest-probability tokens; we approximate $H_{i}$ with the truncated-entropy lower bound, sufficiently accurate when the tail mass is small.
   2. 2.
@@
 In a more automated pipeline, it could function as a triage mechanism for challenging pages.
-5.3 Reviewer Effort
+5.3 Annotator Effort
 Our human annotators reported that spotting error-prone tokens became faster when the three highest-entropy windows were highlighted in color.
@@
 In conclusion, these observations reinforce the general principle that “where the model is least certain” often overlaps with “where real OCR errors lurk.” The sliding-window approach, while simple, appears surprisingly effective as a lightweight mechanism to triage GPT-based OCR outputs for potential mistakes.
+7 Conclusion and Future Work
+In this paper, we presented a novel proof-of-concept framework for localizing OCR errors in GPT-based mathematical transcripts using per-token entropy signals.
+By converting per-token log-probabilities into a practical “uncertainty landscape,” our method pinpointed short contiguous segments—hotspots—most likely to contain recognition problems.
+The sliding-window averaging approach proved to be a lightweight and interpretable means of detecting local failures, as demonstrated on a small-scale corpus of scanned research pages rendered at several resolutions.
+Qualitative evaluation showed that high-entropy windows align with various OCR pitfalls, including missing symbols, mismatched delimiters, and confusions among visually similar characters.
+Although our findings relied on human annotation over a modest set of 36 image samples, the strong anecdotal evidence suggests that this uncertainty-driven technique can help researchers or proofreaders focus attention on precisely those sub-sequences that merit closer review.
+Notably, it operates with minimal overhead, given that most GPT-based APIs already expose top-k token probabilities.
+Nonetheless, we acknowledge certain limitations.
+First, the overall study size remains small, so additional experiments with larger and more diverse datasets are necessary to confirm the generality of our insights.
+Second, our “hotspot” detection occasionally flagged high-entropy regions that did not contain outright errors, pointing to a need for improved thresholds or follow-up diagnostics (e.g., automated re-prompting).
+Finally, the method was tested primarily on text with stationary mathematical expressions—tables, complex layouts, or heavily multi-column documents might require more advanced segmentation and tracking techniques.
+Looking ahead, several extensions and directions emerge:
+  • Automated Re-Prompting. The proposed hotspot detection can be combined with a selective re-querying of GPT, requesting localized corrections only for suspicious subsequences.
+  • Fine-Grained Error Classification.
+Beyond marking high-uncertainty tokens as potential errors, a downstream classifier could distinguish among error types (e.g., missing symbol vs.
+symbol confusion) to inform specialized correction modules.
+  • Adaptive Windowing. Varying the window length automatically in response to local text features (such as formula boundaries) could enhance precision in highlighting exactly the relevant tokens.
+  • Multi-Task and Multilingual Extensions.
+The same sliding-window entropy approach may prove useful in other scenarios (e.g., handwriting recognition, multilingual OCR) where local token competition signals confusion.
+  • Larger-Scale Benchmarking.
+A more expansive and systematic data collection—covering more diverse page layouts, fonts, and scanning artifacts—would facilitate rigorous quantitative comparisons with alternative error-spotting methods.
+In conclusion, our entropy heat-mapping framework illustrates how local confidence signals can offer a powerful lens into GPT-based OCR pipelines.
+By exposing precisely which tokens are most uncertain, it sets the stage for deeper refinements and targeted corrections, thereby improving the reliability of OCR outputs while minimizing human proofreading workloads.
+We hope that this proof-of-concept fosters further research on token-level uncertainty in large language models and drives the development of more reliable, transparent multimodal transcription systems.
+Declarations
+Conflict of Interest
+The authors have no relevant financial or non-financial interests to disclose. The authors have no conflict of interest to declare for the content of this article.
+Code Availability
+The Python code used in this study, developed by the authors, is openly accessible on GitHub at https://github.com/Alexei-WLU/scan2latex-entropy.
+This repository contains all scripts necessary to reproduce the experimental results presented in this paper.